# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ 1 Introduction
Sequential learning algorithms are widely used in practice, even when the entire training sample is known in advance (batch setup), with stochastic gradient descent being one of the most well-known examples. The model we study in this paper represents an intermediate approach, combining sequential procedures with their batch counterparts. A standard sequential regression problem can be described as follows: at round $t$, one observes the design vector $x_{t}$ and makes a prediction $\widehat{y}_{t}$ based on previous observations, aiming to approximate the true value $y_{t}$, which is revealed only after the prediction $\widehat{y}_{t}$ is made. Our setup, often referred to as transductive online learning, provides the learner with additional information, as the entire set of design vectors $x_{1},\ldots,x_{T}$ (though not their order) is revealed to the learner in advance and can be utilized. The term transduction, coined by Vapnik (2013), refers to reasoning directly from specific training examples to specific test examples, unlike induction, which generalizes to general prediction rules.
@@
Our key focus is on the separation between the regret bounds achievable in the transductive setting compared to the standard online setting, particularly in the case of linear regression with squared loss, including sparse classification, as well as classification problems like logistic regression and classification with hinge loss. Some of our key findings are simple enough to be formulated in the introduction. Let us start with classification with logarithmic loss: in the standard online setup (Hazan et al., 2014; Foster et al., 2018), we observe the vectors $x_{t}\in\mathbb{R}^{d}$ round by round, where $t=1,\ldots,T$, and assign the probabilities $\widehat{p}(x_{t},1)$ to label $1$ and $\widehat{p}(x_{t},-1)$ to label $-1$, after which we observe the true label $y_{t}\in\{1,-1\}$. For some $b>0$, our aim is to minimize the regret with respect to the logarithmic loss, that is,
@@
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function. Here, for the sake of example, we assumed $\|x_{t}\|_{2}\leq 1$, though importantly, the results in this paper will completely bypass any boundedness assumptions on the design vectors. The approaches based on choosing the proper predictor, that is, a predictor corresponding to the logistic regression model, $\widehat{p}(x_{t},1)=\sigma(\langle x_{t},\widehat{\theta}_{t}\rangle)$, are quite limited. As shown by Hazan, Koren, and Levy (2014), and rather informally skipping some assumptions, the lower bound holds
@@
It is important to note that the assumption $b<\infty$ in (2) is not natural. For instance, in the favorable scenario where all data points belong to the same class $+1$, the optimal parameter vector $\theta^{\star}$ could have an arbitrarily large norm. As discussed earlier, our focus in this paper is on the transductive setup, where the set of vectors $\{x_{1},\ldots,x_{T}\}$ is known to the learner beforehand. By denoting the regret in this setup as $\operatorname{Reg}_{T}^{\operatorname{trd}}$, we demonstrate that
@@
Additionally, we introduce a variant of the online-to-batch conversion argument to illustrate the application of our results in the statistical setup. In the statistical learning context, one observes i.i.d. sample $(X_{t},Y_{t})_{t=1}^{T}$ of copies of $(X,Y)$, where $X\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$. Our analysis shows that there exists a probability assignment $\widetilde{p}(\cdot,\cdot)$, constructed based on the training sample $(X_{t},Y_{t})_{t=1}^{T}$, such that
@@
with expectations taken over both the training sample $(X_{t},Y_{t})_{t=1}^{T}$ and an independent test observation $(X,Y)$. A key feature of our excess risk bound is that the right-hand side of (4) is completely independent of the distributions of $X$ and $Y$. Moreover, the curvature of the logarithmic loss function enables the $d\log(T)/T$ convergence rate without requiring additional realizability/low noise assumptions. Despite the extensive literature on logistic regression in the statistical framework (for comparison, see, e.g., (Bach, 2010; Hazan et al., 2014; Plan et al., 2017; Ostrovskii and Bach, 2021; Puchkin and Zhivotovskiy, 2023; Kuchelmeister and van de Geer, 2024; Hsu and Mazumdar, 2024) for a non-exhaustive list of existing risk upper bounds in this context), this fully assumption-free convergence rate has not been previously established even for mixtures of sigmoids proposed in (Kakade and Ng, 2004; Foster et al., 2018) and other batch-specific results as in (Mourtada and Gaïffas, 2022; Vijaykumar, 2021; van der Hoeven et al., 2023). We emphasize again that our approach relies on using variants of the continuous exponential weights algorithm.
Our findings go well beyond the logarithmic loss. For example, we consider the case of sparse linear regression with squared loss. Specifically, in the transductive setting with squared loss, assuming $|y_{t}|\leq m$ for all $t=1,\ldots,T$ and comparing against any $\theta^{\star}$ with $\|\theta^{\star}\|_{0}\leq s$, we prove a regret bound for a computationally efficient method:
@@
where the dependence on the compatibility constant $\kappa_{s}$ is logarithmic, contrasting with the polynomial dependence found in most known results for computationally efficient predictors that satisfy sparsity regret or risk bounds.
Furthermore, in the setup of classification with hinge loss, we prove the so-called first-order regret bounds. In more detail, we observe the vectors $x_{t}\in\mathbb{R}^{d}$ round by round, where $t=1,\ldots,T$, and assign the real-valued output to label 1 and $\widehat{f}(x_{t})$, trying to minimize, for some $\gamma>0$, the total hinge loss:
@@
In the transductive setup, where we know the set $\{x_{1},\ldots,x_{T}\}$, if additionally the data is linearly separable with margin $\gamma$, the result of our Theorem 4.1 implies, in particular, that
@@
Prediction is again achieved using a version of the continuous exponential weights algorithm, and the right-hand side of the above cumulative error bound does not depend on $\gamma$, which must be the case when the set $\{x_{1},\ldots,x_{T}\}$ is unknown.
@@ 1.1 Key contributions and structure of the paper
Our key contributions can be summarized through the regret bounds we derive for the transductive online learning framework, considering the squared loss, logistic loss, and hinge loss. These bounds explicitly reveal distinctions between the transductive and inductive setups across different loss functions. Moreover, our algorithm exhibits computational advantages through sampling from log-concave densities. Concretely, we summarize them as the following:
@@ 1. $\bullet$
    In Section 2, we revisit the classical problem of online squared loss regression. We recover the optimal rate of $O(dm^{2}\log T)$ for the transductive setup originally established by Gaillard et al. (2019) with a transductive prior without making any assumptions about the boundedness of covariates or the norm of the optimal solution (comparator). In the sparse linear regression setting, we introduce a sparse prior that achieves assumption-free bounds with efficient computation via sampling from log-concave densities. This computation (detailed in Section 5) relies on the predictor induced by mixability, expressed as the log ratio of normalizing constants for log-concave densities. We thus exploit a natural connection between the mixability of the squared loss and the recent line of work (Brosse et al., 2018; Ge et al., 2020) on the efficient computations of normalizing constants of log-concave densities. In fact, our result solves a question on computationally efficient sparse online linear regression raised in the monograph (Lattimore and Szepesvári, 2020, Chapter 23). For completeness, we also include results by combining our transductive prior and the discrete sparse prior proposed by Rigollet and Tsybakov (2011, 2012) in Appendix A, where a better regret bound for sparse linear regression is achieved while not being always computationally efficient.
@@ 3. $\bullet$
    In Section 4, we investigate the hinge loss in both the inductive and transductive setup. The hinge loss is unbounded and not mixable. Nevertheless, we present the first-order regret bound scaling as $O(d\log(bT/\gamma))$ in the realizable case, where $\gamma$ is the margin of separation in the inductive case. This regret bound is the parametric counterpart of the celebrated $O(1/\gamma^{2})$ mistake bound for perceptron. We then combine the slab-expert idea with the aforementioned clipping trick to obtain an assumption-free first-order regret bound of $O(d\log T)$ in the transductive setup. The separation between the inductive and transductive setups also exists for the hinge loss.
@@ 4. $\bullet$
    In Section 5, we highlight our computational benefit through sampling from log-concave densities. More concretely, since the squared loss and the logistic loss are convex, the exponential weights algorithm, therefore, predicts log-concave densities. We showcase how to compute the exponential weights algorithm for sparse linear regression and logistic regression. For the latter, we demonstrate how to sample the slab-experts.
@@ 2 Linear Regression with Transductive Priors
Theorem 2.1 (Gaillard, Gerchinovitz, Huard, and Stoltz (2019)).
Assume that the sequence $\{(x_{t},y_{t})\}_{t=1}^{T}$, where $(x_{t},y_{t})\in\mathbb{R}^{d}\times\mathbb{R}$ for all $1\leq t\leq T$, is such that $\max_{t}|y_{t}|\leq m$ and that the matrix $\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$ is invertible [^1] This assumption can be easily bypassed. One way to avoid this is by restricting the problem to the linear span of $x_{1},\ldots,x_{T}$.. For $x\in\mathbb{R}^{d}$, consider the sequence of estimators
@@
The proof of Theorem 2.1 in (Gaillard et al., 2019) employs a linear transformation of the feature vectors, thereby reducing the problem to the classical regret bounds of Vovk (2001) and Azoury and Warmuth (2001) using linear algebraic arguments. Our approach, instead, follows the original path of Vovk (2001) by exploiting the mixability of the loss (details to follow), with the main difference being the choice of a prior distribution that depends on the design vectors $x_{1},\ldots,x_{T}$. Theorem 2.1 can be immediately recovered from the general bounds for the exponential weights algorithm and the mixability of the squared loss discussed in (Vovk, 2001) by choosing the following multivariate Gaussian prior $\mu$ in $\mathbb{R}^{d}$:
@@
We do not provide the derivation here but instead present several versions of the derivation throughout the paper. This direct approach is a natural first step when analyzing the logarithmic loss, hinge loss, and absolute loss. In Appendix C, we discuss some general deterministic identities for online linear regression that are capable of recovering existing sequential linear regression bounds, including the regret bound of Theorem 2.1.
The form of the prior (6) has a close connection with the Jeffreys prior, where $\sqrt{\det\left(\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}\right)}$ plays the role of this prior and the term $\exp\left(-\lambda\theta^{\top}\left(\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}\right)\theta\right)$ plays the role of the exponential tilting that makes this prior integrable. We refer to (Barron, 1999, Section 3) as well as (Polyanskiy and Wu, 2024, Chapter 13.4). Connections with Jeffreys priors in the context of existing information theory results will be made throughout the paper.
@@ 2.1 Sparse linear regression
In this section, as an initial demonstration of our approach utilizing transductive priors for exponential weights, we aim to extend Theorem 2.1. Specifically, we will focus on linear regression, with an emphasis on deriving sparsity regret bounds within the transductive setup. We remark that the techniques used in this section can be readily applied in other contexts, including logistic regression and classification with hinge loss.
@@
The prior is similar to the sparse priors used in (Dalalyan and Tsybakov, 2008, 2012a, 2012b) and (Gerchinovitz, 2013), but, as will always be the case in this paper, it explicitly depends on $x_{1},\ldots,x_{T}$. We now need to define the estimator, which has an explicit integral form. Given $m$, which is the upper bound on $\max_{t}|y_{t}|$, define for $x\in\mathbb{R}^{d}$,
[…]
@@
Definition 2.1 (Smallest scaled singular value).
@@
Theorem 2.2.
Assume that the sequence $\{(x_{t},y_{t})\}_{t=1}^{T}$, where $x_{t},y_{t}\in\mathbb{R}^{d}\times\mathbb{R}$ for all $1\leq t\leq T$, is such that $\max_{t}|y_{t}|\leq m$ and that the smallest scaled singular value condition is satisfied with constant $\kappa_{s}$. Then, for any $s$-sparse $\theta^{\star}\in\mathbb{R}^{d}$, the following regret bound holds for the estimator (8) with $\tau=m\sqrt{\frac{s}{d}}$,
@@
It is time to introduce the exponential weights algorithm (Littlestone and Warmuth, 1994; Vovk, 1990; Freund and Schapire, 1997). Fix $\eta>0$ and let $\ell_{\theta}(\cdot):\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a set of loss functions parameterized by some $\Theta\subseteq\mathbb{R}^{d}$. For the notation simplicity, we usually use the same symbols for distributions and their densities. Fix some prior $\mu$ over $\Theta$. Given the sequence $(x_{1},y_{1}),\ldots,(x_{T},y_{T})$, define $\rho_{1}=\mu$. For $t\geq 2$, we set
@@
Lemma 2.2.
In the setup of Lemma 2.1, the following identity holds:
$-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\nolimits_{i=1}^{T}\eta\ell_{\theta}(x_{t},y_{t})\right)\right)=\inf_{\varphi}\left(\mathbb{E}_{\theta\sim\varphi}\sum\limits_{t=1}^{T}\ell_{\theta}(x_{t},y_{t})+\frac{\mathcal{KL}(\varphi\|\mu)}{\eta}\right).$
@@
Lemma 2.3 (Mixability of the squared loss).
@@
Proof of Theorem 2.2 In what follows, $\eta=\frac{1}{2m^{2}}$. Our first observation is that if we choose $\rho=\rho_{t}$ in Lemma 2.3, where $\rho_{t}$ is the exponential weights algorithm with prior $\mu$ given by (7) and $\ell_{\theta}(x,y)=(y-\langle x,\theta\rangle)^{2}$, then the predictor (12) coincides exactly with the predictor $\widehat{f}_{t}$ given by (8). This implies by Lemma 2.3, Lemma 2.1 and Lemma 2.2:
$\sum\limits_{t=1}^{T}(y_{t}-\widehat{f}_{t}(x_{t}))^{2}$ $\leq\sum\limits_{t=1}^{T}-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\exp\left(-\eta(y_{t}-\langle x_{t},\theta\rangle))^{2}\right)\right)$
$=-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\nolimits_{t=1}^{T}\eta(y_{t}-\langle x_{t},\theta\rangle))^{2}\right)\right)$
@@
By symmetry we have $\mathbb{E}_{\theta\sim\mu^{\star}}\theta=\theta^{\star}$. Moreover, a simple computation shows that $\mathbb{E}_{\theta\sim\mu}(\theta^{(j)})^{2}=\tau^{2}/(\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2})$. Therefore, we have
@@
We make several remarks regarding the result. The bound of Theorem 2.2 depends logarithmically on $\frac{1}{\kappa_{s}}$. However, even if $\kappa_{s}$ is close to zero, our bound remains meaningful as it depends at most logarithmically on $\|\theta^{\star}\|$ and $\max_{t}\|x_{t}\|_{2}$, provided both are bounded. This follows directly from upper bounding (14) when both $\theta^{\star}$ and all $x_{t}$-s are bounded.

### v3

@@ Abstract
Keywords: Transductive online learning, unbounded losses, logistic regression, hinge loss.
@@ 1 Introduction
Sequential learning algorithms are widely used in practice, even when the entire training sample is known in advance (batch setup), with stochastic gradient descent being one of the most well-known examples. The model we study in this paper represents an intermediate approach, combining sequential procedures with their batch counterparts. A standard sequential regression problem can be described as follows: at round $t$, one observes the design vector $x_{t}$ and makes a prediction $\widehat{y}_{t}$ based on previous observations, aiming to approximate the true value $y_{t}$, which is revealed only after the prediction $\widehat{y}_{t}$ is made. Our setup, often referred to as transductive online learning, provides the learner with additional information, as the entire set of design vectors $x_{1},\ldots,x_{T}$ (though not their order) is revealed to the learner in advance and can be utilized. In the terminology of Vapnik (2013), transduction refers to situations where the aim is not to infer a general prediction rule for arbitrary future points, but to predict the labels of a specific, fixed set of points known beforehand. This exactly matches our setting, where the goal is to assign labels to a predetermined sequence rather than to predict each point in an arbitrary sequence as it arrives.
@@
Our key focus is on the separation between the regret bounds achievable in the transductive setting compared to the standard online setting, particularly in the case of linear regression with squared loss, including sparse linear regression, as well as classification problems like logistic regression and classification with hinge loss. Some of our key findings are simple enough to be formulated in the introduction. Let us start with classification with logarithmic loss: in the standard online setup (Hazan et al., 2014; Foster et al., 2018), we observe the vectors $x_{t}\in\mathbb{R}^{d}$ round by round, where $t=1,\ldots,T$, and assign the probabilities $\widehat{p}(x_{t},1)$ to label $1$ and $\widehat{p}(x_{t},-1)$ to label 1-1, after which we observe the true label $y_{t}\in\{1,-1\}$. For some $b>0$, our aim is to minimize the regret with respect to the logarithmic loss, that is,
@@
where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function. Here, for the sake of example, we assumed $\|x_{t}\|_{2}\leq 1$, though importantly, the results in this paper will completely bypass any boundedness assumptions on the design vectors. The approaches based on choosing the proper predictor, that is, a predictor corresponding to the logistic regression model, $\widehat{p}(x_{t},1)=\sigma(\langle x_{t},\widehat{\theta}_{t}\rangle)$, where $\widehat{\theta}_{t}$ is independent of $x_{t}$, are quite limited. As shown by Hazan, Koren, and Levy (2014), and rather informally skipping some assumptions, the following lower bound holds:
@@
It is important to note that the assumption $b<\infty$ in (2) is not natural. For instance, in the favorable scenario where all data points belong to the same class +1+1, the optimal parameter vector $\theta^{\star}$ could have an arbitrarily large norm. As discussed earlier, our focus in this paper is on the transductive setup, where the set of vectors $\{x_{1},\ldots,x_{T}\}$ is known to the learner beforehand. By denoting the regret in this setup as $\operatorname{Reg}_{T}^{\operatorname{trd}}$, we demonstrate that
@@
[…]
@@
with expectations taken over both the training sample $(X_{t},Y_{t})_{t=1}^{T}$ and an independent test observation $(X,Y)$. A key feature of our excess risk bound is that the right-hand side of (4) is completely independent of the distributions of $X$ and $Y$. Moreover, the curvature of the logarithmic loss function enables the $d\log(T)/T$ convergence rate without requiring additional realizability/low noise assumptions. Despite the extensive literature on logistic regression in the statistical framework (for comparison, see, e.g., (Bach, 2010; Hazan et al., 2014; Plan et al., 2017; Ostrovskii and Bach, 2021; Puchkin and Zhivotovskiy, 2023; Kuchelmeister and van de Geer, 2024; Hsu and Mazumdar, 2024) for a non-exhaustive list of existing risk upper bounds in this context), this fully assumption-free convergence rate has not been previously established even for mixtures of sigmoids proposed in (Kakade and Ng, 2004; Foster et al., 2018) and other batch-specific results as in (Mourtada and Ga¨ıffas, 2022; Vijaykumar, 2021; van der Hoeven et al., 2023). We emphasize again that our approach relies on using variants of the continuous exponential weights algorithm.
Our findings go well beyond the logarithmic loss. For example, we consider the case of sparse linear regression with squared loss. Specifically, in the transductive setting with squared loss, assuming $|y_{t}|\leq m$ for all $t=1,\ldots,T$ and comparing against any $\theta^{\star}$ with $\|\theta^{\star}\|_{0}\leq s$, we prove a regret bound of
@@
where the dependence on the compatibility constant $\kappa_{s}$ is logarithmic, contrasting with the polynomial dependence found in most known results for predictors that satisfy sparsity regret or risk bounds.
Furthermore, in the setup of classification with hinge loss, we prove the so-called first-order regret bounds. In more detail, we observe the vectors $x_{t}\in\mathbb{R}^{d}$ round by round, where $t=1,\ldots,T$, and output a real-valued score $\widehat{f}(x_{t})$, trying to minimize, for some $\gamma>0$, the total hinge loss:
@@
In the transductive setup, where we know the set $\{x_{1},\ldots,x_{T}\}$, if additionally the data is linearly separable with margin $\gamma$, the result of our Theorem 21 implies, in particular, that
@@
Prediction is again achieved using a version of the continuous exponential weights algorithm, and the right-hand side of the above cumulative error bound does not depend on $\gamma$, which need not hold in the inductive (unknown-set) setup.
@@ 1.1 Key contributions and structure of the paper
Our key contributions can be summarized through the regret bounds we derive for the transductive online learning framework, considering the squared loss, logistic loss, and hinge loss. These bounds explicitly reveal distinctions between the transductive and inductive setups across different loss functions. Moreover, our algorithm exhibits computational advantages through sampling from log-concave densities. Concretely, we summarize them as follows:
@@ 1. $\bullet$
    In Section 2, we revisit the classical problem of online squared loss regression. We recover the optimal rate of $O(dm^{2}\log T)$ for the transductive setup originally established by Gaillard et al. (2019) with a transductive prior without making any assumptions about the boundedness of covariates or the norm of the optimal solution (comparator). In the sparse linear regression setting, we introduce a sparse prior that achieves regret bounds independent of the norm of $\theta^{\star}$ via the evaluation of Gaussian-type integrals. For completeness, we also include results by combining our transductive prior and the discrete sparse prior proposed by Rigollet and Tsybakov (2011, 2012) in Appendix A, where a better regret bound for sparse linear regression is achieved, albeit through a method involving sampling from discrete measures, which may require a prohibitively large number of samples from these distributions in the worst case.
@@ 3. $\bullet$
    In Section 4, we investigate the hinge loss in both the inductive and transductive setup. The hinge loss is unbounded and not mixable. Nevertheless, we present the first-order regret bound scaling as $O(d\log(bT/\gamma))$ in the realizable case, where $\gamma$ is the margin of separation in the inductive case and $b$ is the norm of the optimal weight vector. This regret bound is the parametric counterpart of the celebrated $O(1/\gamma^{2})$ mistake bound for perceptron. We then combine the slab-expert idea with the aforementioned clipping trick to obtain an assumption-free first-order regret bound of $O(d\log T)$ in the transductive setup. The separation between the inductive and transductive setups also exists for the hinge loss.
@@ 4. $\bullet$
    In Section 5, we highlight our computational benefit through sampling from log-concave densities. More concretely, since the squared loss and the logistic loss are convex, the exponential weights algorithm, therefore, predicts log-concave densities. We showcase how to compute the exponential weights algorithm for sparse linear regression and logistic regression. For the latter, we demonstrate how to sample the slab-experts, though our constructions lead to computational complexities exponential in dimension in the worst case.
@@ 2 Linear Regression with Transductive Priors
Theorem 1 (Gaillard, Gerchinovitz, Huard, and Stoltz (2019)).
Assume that the sequence $\{(x_{t},y_{t})\}_{t=1}^{T}$, where $(x_{t},y_{t})\in\mathbb{R}^{d}\times\mathbb{R}$ for all $1\leq t\leq T$, is such that $\max_{t}|y_{t}|\leq m$ and that the matrix $\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}$ is invertible [^1] This assumption can be easily bypassed. One way around this is by restricting the problem to the linear span of $x_{1},\ldots,x_{T}$.. For $x\in\mathbb{R}^{d}$, consider the sequence of estimators
@@
The proof of Theorem 1 in (Gaillard et al., 2019) employs a linear transformation of the feature vectors, thereby reducing the problem to the classical regret bounds of Vovk (2001) and Azoury and Warmuth (2001) using linear algebraic arguments. Our approach, instead, follows the original path of Vovk (2001) by exploiting the mixability of the loss (details to follow), with the main difference being the choice of a prior distribution that depends on the design vectors $x_{1},\ldots,x_{T}$. Theorem 1 can be immediately recovered from the general bounds for the exponential weights algorithm and the mixability of the squared loss discussed in (Vovk, 2001) by choosing the following multivariate Gaussian prior $\mu$ in $\mathbb{R}^{d}$:
@@
We do not provide the derivation here but instead present several versions of the derivation throughout the paper. This direct approach is a natural first step when analyzing the logarithmic loss, hinge loss, and absolute loss. In Appendix C, we discuss some general deterministic identities for online linear regression that are capable of recovering existing sequential linear regression bounds, including the regret bound of Theorem 1.
The prior (6) has a close connection with the Jeffreys prior, where $\sqrt{\det\left(\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}\right)}$ plays the role of this prior and the term $\exp\left(-\lambda\theta^{\top}\left(\sum\nolimits_{t=1}^{T}x_{t}x_{t}^{\top}\right)\theta\right)$ plays the role of the exponential tilting that makes this prior integrable. We refer to (Barron, 1999, Section 3) as well as (Polyanskiy and Wu, 2024, Chapter 13.4). Connections with Jeffreys priors in the context of existing information theory results will be made throughout the paper.
@@ 2.1 Sparse linear regression
In this section, as an initial demonstration of our approach utilizing transductive priors for exponential weights, we aim to extend Theorem 1. Specifically, we will focus on linear regression, with an emphasis on deriving sparsity regret bounds within the transductive setup. We remark that the techniques used in this section can be readily applied in other contexts, including logistic regression and classification with hinge loss.
@@
The prior is similar to the sparse priors used in (Dalalyan and Tsybakov, 2008, 2012a, 2012b) and (Gerchinovitz, 2013), but, as will always be the case in this paper, it explicitly depends on $x_{1},\ldots,x_{T}$. The estimator we use has an explicit integral form following Vovk’s algorithm (Lemma 6). Given $m$, which is the upper bound on $\max_{t}|y_{t}|$, define for $x\in\mathbb{R}^{d}$,
$\widehat{f}_{t}(x)=\frac{m}{2}\cdot\log\left(\frac{\int\limits_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2m^{2}}(m-\langle x,\theta\rangle)^{2}-\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}\right)\mu(\theta)d\theta}{\int\limits_{\mathbb{R}^{d}}\exp\left(-\frac{1}{2m^{2}}(-m-\langle x,\theta\rangle)^{2}-\frac{1}{2m^{2}}\sum\limits_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}\right)\mu(\theta)d\theta}\right).$ (8)
@@
Definition 2 (Smallest scaled singular value).
@@
Theorem 3.
Assume that the sequence $\{(x_{t},y_{t})\}_{t=1}^{T}$, where $(x_{t},y_{t})\in\mathbb{R}^{d}\times\mathbb{R}$ for all $1\leq t\leq T$, is such that $\max_{t}|y_{t}|\leq m$ and that the smallest scaled singular value condition is satisfied with constant $\kappa_{s}$. Then, for any $s$-sparse $\theta^{\star}\in\mathbb{R}^{d}$, the following regret bound holds for the estimator (8) with $\tau=m\sqrt{\frac{s}{d}}$,
@@
It is time to introduce the exponential weights algorithm (Littlestone and Warmuth, 1994; Vovk, 1990; Freund and Schapire, 1997). Fix $\eta>0$ and let $\ell_{\theta}(\cdot):\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a set of loss functions parameterized by some $\Theta\subseteq\mathbb{R}^{d}$. For notational simplicity, we usually use the same symbols for distributions and their densities. Fix some prior $\mu$ over $\Theta$. Given the sequence $(x_{1},y_{1}),\ldots,(x_{T},y_{T})$, define $\rho_{1}=\mu$. For $t\geq 2$, we set
@@
Lemma 5.
In the setup of Lemma 4, the following identity holds:
$\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\nolimits_{t=1}^{T}\eta\ell_{\theta}(x_{t},y_{t})\right)\right)=\inf_{\varphi}\left(\mathbb{E}_{\theta\sim\varphi}\sum\limits_{t=1}^{T}\ell_{\theta}(x_{t},y_{t})+\frac{\mathcal{KL}(\varphi\|\mu)}{\eta}\right).$
@@
Lemma 6 (Mixability of the squared loss).
@@
Proof of Theorem 3 In what follows, $\eta=\frac{1}{2m^{2}}$. Our first observation is that if we choose $\rho=\rho_{t}$ in Lemma 6, where $\rho_{t}$ is the exponential weights algorithm with prior $\mu$ given by (7) and $\ell_{\theta}(x,y)=(y-\langle x,\theta\rangle)^{2}$, then the predictor (12) coincides exactly with the predictor $\widehat{f}_{t}$ given by (8). This implies by Lemma 6, Lemma 4 and Lemma 5:
$\sum\limits_{t=1}^{T}(y_{t}-\widehat{f}_{t}(x_{t}))^{2}$ $\leq\sum\limits_{t=1}^{T}-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\exp\left(-\eta(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right)$
$=-\frac{1}{\eta}\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\nolimits_{t=1}^{T}\eta(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right)$
@@
By symmetry $\mathbb{E}_{\theta\sim\mu^{\star}}\theta=\theta^{\star}$. Moreover, a simple computation shows that $\mathbb{E}_{\theta\sim\mu}(\theta^{(j)})^{2}=\tau^{2}/(\sum\nolimits_{t=1}^{T}(x_{t}^{(j)})^{2})$ (recall that $\mu$ is defined by (7)). Therefore, we have
@@
We make several remarks regarding the result. The bound of Theorem 3 depends logarithmically on $\frac{1}{\kappa_{s}}$. However, even if $\kappa_{s}$ is close to zero, our bound remains meaningful as it depends at most logarithmically on $\|\theta^{\star}\|$ and $\max_{t}\|x_{t}\|_{2}$, provided both are bounded. This follows directly from upper bounding (14) when both $\theta^{\star}$ and all $x_{t}$-s are bounded.